OpenStudy (nuccioreggie):

What is the value of the expression All of 4.8 multiplied by 10 to the power 9 over all of 1.6 multiplied by 10 to the power 3? 3.0 × 106 3.0 × 1012 3.2 × 108 3.2 × 1012

OpenStudy (michele_laino):

what is 4.8 * 1.6=...?

OpenStudy (nuccioreggie):

I think its B

OpenStudy (michele_laino):

sorry... what is: 4.8/1.6=...?

OpenStudy (michele_laino):

hint: \[\Large \frac{{{{10}^9}}}{{{{10}^6}}} = {10^{\left( {9 - 6} \right)}} = {10^{\left( {...} \right)}}\]

OpenStudy (nuccioreggie):

3

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (michele_laino):

sorry... \[\Large \frac{{{{10}^9}}}{{{{10}^3}}} = {10^{\left( {9 - 3} \right)}} = {10^{\left( {...} \right)}}\]

OpenStudy (nuccioreggie):

A

OpenStudy (michele_laino):

that's right!